OpenStudy (anonymous):

1000 (X) 12 Divide 2 = Don't Feel Like Finding My Calculate So Can Yall Help Me

OpenStudy (unklerhaukus):

\[1000\times\frac{12} 2\]

OpenStudy (anonymous):

Whats The Answer Lol

OpenStudy (unklerhaukus):

\[\frac{2\times500\times12} 2\] \[=\frac{\cancel 2\times500\times12}{\cancel 2}\]

OpenStudy (unklerhaukus):

\[500\times12=500\times10+500\times2\]

OpenStudy (anonymous):

which is equal to 6000

OpenStudy (anonymous):

Well, multiplying 10 by something is easy right? You just add the zero to the end? My biggest hint would be to first do the math for \[12 \over 6\] and then multiply that by 1000.

OpenStudy (anonymous):

Thanks You Guys ((:

OpenStudy (lgbasallote):

@mzPeaches if you just need a calculator then use http://wolframalpha.com it's so much smarter to use it wherein you'll get an accurate answer rather than bothering yourself here waiting for an answer isnt it?